# Qizhe Zhang1,2  Aosong Cheng1  Ming Lu3  Zhiyong Zhuo1  Minqi Wang1 Jiajun Cao1  Shaobo Guo2  Qi She2  Shanghang Zhang1

### v1

@@ Abstract
Large vision-language models (VLMs) often rely on a substantial number of visual tokens when interacting with large language models (LLMs), which has proven to be inefficient. Recent efforts have aimed to accelerate VLM inference by pruning visual tokens. Most existing methods assess the importance of visual tokens based on the text-visual cross-attentions in LLMs. In this study, we find that the cross-attentions between text and visual tokens in LLMs are inaccurate. Pruning tokens based on these inaccurate attentions leads to significant performance degradation, especially at high reduction ratios. To this end, we introduce FasterVLM, a simple yet effective training-free visual token pruning method that evaluates the importance of visual tokens more accurately by utilizing attentions between the [CLS] token and image tokens from the visual encoder. Since FasterVLM eliminates redundant visual tokens immediately after the visual encoder, ensuring they do not interact with LLMs and resulting in faster VLM inference. It is worth noting that, benefiting from the accuracy of [CLS] cross-attentions, FasterVLM can prune 95% of visual tokens while maintaining 90% of the performance of LLaVA-1.5-7B. We apply FasterVLM to various VLMs, including LLaVA-1.5, LLaVA-NeXT, and Video-LLaVA, to demonstrate its effectiveness. Experimental results show that our FasterVLM maintains strong performance across various VLM architectures and reduction ratios, significantly outperforming existing text-visual attention-based methods. Our code is available at https://github.com/Theia-4869/FasterVLM.
@@ 1 Introduction
Figure 1: Comparison of LLaVA-1.5-7B on a range of 10 benchmarks, with a 90% reduction ratio. FasterVLM outperforms other methods by a large margin on 9 tasks and achieves comparable results on the remaining one.
Figure 2: Relationship between performance and efficiency of different methods for LLaVA-1.5-7B. The $x$-axis stands for the reduction ratio of visual tokens, and the $y$-axis stands for the averaged performance of 10 benchmarks. While other methods suffer significant performance degradation as the token reduction ratio increases, FasterVLM maintains most performance (90% of vanilla LLaVA-1.5-7B) even at a high reduction ratio (95%).
@@
With recent advances in large language models (LLMs) [44, 3, 43, 64, 54, 55, 13, 38], numerous efforts have been made to extend their powerful reasoning capabilities to multi-modal tasks [14, 35, 62], giving rise to vision language models (VLMs) such as GPT-4 [1], Gemini [51], LLaVA [34], and Qwen-VL [2]. These VLMs typically contain a visual encoder (e.g. CLIP) to convert visual inputs into sequential representations, and an LLM (e.g. Llama 2) for text generation, with modal alignment for the LLM to accept visual tokens as input [34]. Although these approaches have achieved remarkable performance in multi-modal tasks like visual question answering [15, 20, 16, 36, 50], the visual tokens, which comprise the majority of the input sequence, significantly increase the computational complexity and cost of VLM inference. Some work improves model performance by increasing input image resolution [37, 33, 59], but this leads to longer visual token sequences, further escalating computational overhead, which can be even worse in video models like Video-LLaVA [30].
To reduce the inference cost of VLMs, FastV [8] analyses the information flow in LLMs within VLMs and identifies the inefficient visual attention phenomenon: In the deeper layers of the LLM (after layer 2), visual tokens receive significantly less attention than their textual counterparts, implying that less useful information flows from visual tokens to the text portion, motivating the pruning of visual tokens in VLMs. According to the distribution of attention in different layers, FastV proposes an intuitive solution, ranking the visual tokens based on attention scores received after layer 2 of the LLM and pruning the lowest $R\%$ according to the computational budget, thereby reducing visual token redundancy and improving VLM inference efficiency. Subsequently, more work follows this paradigm [61, 65], utilizing various methods to prune visual tokens through text-visual attention from the LLM. Although FastV successfully identifies the inefficiency of visual tokens in VLMs, pruning based on text-visual attention is not effective. As shown in Fig. 2, the performance of VLM degrades significantly as the reduction ratio increases.
In this paper, we find that the text-visual attention within the LLMs of current VLMs fails to align well with the actual significance of visual tokens, meaning that the attention weights assigned to each visual token do not correspond appropriately to their importance in the task. This misalignment can be explained from two aspects, termed attention shift and attention dispersion. Specifically, attention shift refers to a tendency for textual attention to focus more on later parts of the visual token sequence, which is not desirable for preserving valuable visual information. We attribute this phenomenon to two main factors: the unidirectional information flow within the LLM and the relative positional relationships between tokens. These factors cause textual prompts to preferentially focus on the visual tokens positioned later in the sequence. Conversely, Attention dispersion refers to the less concentrated attention distribution within the LLM compared to the visual encoder. In the LLM, more visual tokens receive relatively high attention scores, but the highest attention value is lower. We believe that, despite the multi-modal projector’s role in aligning visual representations with text embeddings, a gap still remains between them. This gap results in noisy text-visual attention at shallower layers, hindering the effective pruning of important visual tokens.
To more accurately evaluate the importance of visual tokens, we introduce FasterVLM, a simple yet effective pruning method. FasterVLM directly utilizes the attention weights from the [CLS] token within the image encoder to assess the importance of each visual token. By removing tokens with lower attention scores before they are processed by the LLM, we achieve maximum inference acceleration. Quantitative and qualitative analyses indicate that [CLS] attention does not exhibit the shift phenomenon, thanks to the global attention mechanism in the image encoder. Moreover, compared to text-visual attention, [CLS] attention is more concentrated, serving as a better indicator for evaluating the importance of visual tokens. As noted in [11], certain artifacts exist in the attention maps of Vision Transformers [53, 45, 42], where the regions contain pixels highly similar to their surroundings. The model discards local information from these regions and stores global information in these artifacts. We find that these global tokens significantly impact VLM performance. Under high reduction ratios, previous methods struggle to identify these artifacts due to inaccurate text-visual attention. In contrast, FasterVLM can accurately identify these visual tokens containing global information via [CLS] attention, preserving most of the performance.
@@
Benefiting from the simple design, FasterVLM can be easily applied to different VLMs, including LLaVA-1.5 [32], LLaVA-NeXT [33], and Video-LLaVA [30]. As shown in Fig. 2, extensive experiments on various vision-language benchmarks demonstrate that FasterVLM consistently outperforms other text-visual attention-based methods under different reduction ratios. For example, FasterVLM is capable of pruning 95% of visual tokens in LLaVA-1.5-7B, reducing inference FLOPs by more than 95%, and maintaining 90% of the original performance across 10 tasks without requiring any additional training (details for each task are provided in Fig. 2). In scenarios involving longer visual token sequences—such as high-resolution images (LLaVA-NeXT) or video inputs (Video-LLaVA)—FasterVLM continues to achieve superior performance with its simple design.
In summary, the contributions of our work are threefold:
@@
    We uncover that the inaccuracies of text-visual attention in the LLM—stemming from attention shift and attention dispersion—compromise its effectiveness for visual token pruning, and we provide a detailed explanation of these phenomena.
@@
    We introduce FasterVLM, a simple yet effective training-free visual token pruning method that evaluates the importance of visual tokens more accurately by the cross-attentions between the [CLS] token and image tokens from the visual encoder.
@@
    We apply FasterVLM to different types of VLMs and conduct extensive experiments on various vision-language tasks, showing that FasterVLM consistently outperforms existing methods and maintains high performance even under extreme reduction ratios.
@@ 2.1 Vision-Language Models (VLMs)
The recent impressive success of large language models (LLMs) [44, 3, 43, 64, 54, 55, 13, 38] leads to a trend of extending their powerful reasoning capabilities to multi-modal comprehension tasks, giving birth to vision-language models (VLMs) [1, 51]. These VLMs generally include a visual encoder for serializing representations of input images and an LLM for text generation. To make the LLM acceptable for visual representations as input, VLMs typically align the visual and language modalities with an alignment module, which can be a simple linear layer [34], an MLP projector [32], or a deep query-based network [26, 2]. Although this allows the LLM to have visual perception, the introduction of long visual token sequences increases the computational burdens. Furthermore, studies have shown that existing VLMs still suffer from visual shortcomings [52] or hallucinations [18]. To mitigate this, efforts have been made to improve VLM performance by increasing the resolution of input images [37, 59], which further exacerbates the computation. For example, LLaVA-1.5 [34] encodes an image with 336 resolution into 576 visual tokens, while LLaVA-NeXT [33] doubles the resolution, resulting in 2880 tokens, potentially dozens of times more than the textual prompts. For video inputs, Video-LLaVA [30] needs to handle even more visual tokens from multiple frames, making VLM inference prohibitively expensive. Optimizing the inference process of VLMs is an urgent task to enable their application in resource-constrained real-world scenarios.
2.2 Token Compression for VLMs
One way to optimize VLM inference is by compressing the visual tokens that occupy the majority of the input sequence. Several studies have explored token sequence compression in language models [9, 19, 41, 46]. Compared with text, image information tends to have higher redundancy, making visual token compression for VLMs more reasonable and effective. LLaVA-PurMerge [48] leverages attention mechanisms to select important visual tokens and merges them using similar key clustering, achieving competitive performance while improving VLM inference efficiency. LLaVolta [7] proposes a heuristic and stage-wise compression method that reduces VLM training costs while largely maintaining original performance. TokenPacker [27] adopts a coarse-to-fine approach, using point-to-region attention to retain richer details while compressing visual tokens. MQT [17] and M3 [5] employ Matryoshka Representation Learning [25] to hierarchically compress visual tokens, reducing the number of tokens to no more than 10. However, these methods require additional training for VLMs to adapt to the compressed visual representations. In this work, we focus on VLM inference optimization without additional training.
2.3 Token Pruning for VLMs
Token pruning is another approach to improve model inference efficiency by removing less important tokens from the sequence. With the emergence of Transformer-based networks, this technique has been actively explored in natural language processing [60, 22] and computer vision [47, 40]. However, effective pruning of visual tokens in VLMs still remains under-explored. FastV [8] first identifies the redundancy and inefficiency of visual tokens in VLMs and proposed a simple method to remove visual tokens with low attention scores after the layer 2 of LLM. Subsequently, FitPrune [61] introduces a method to fit pruning recipes based on attention statistics, while SparseVLM [65] removes distractions from text prompts and uses more accurate text attention to progressively sparsify visual tokens. However, these approaches rely on text-visual attention within the LLM to evaluate the importance of visual tokens. In this work, we show that such attention does not align well with visual token importance and propose that pruning based solely on the [CLS] attention from the visual encoder yields unexpectedly good performances.
3 Inaccurate Text-Visual Attention in VLMs
Text-visual attention from the LLM Decoder in VLMs is often used to evaluate the importance of visual tokens for pruning. However, we find this attention does not align with the actual importance of visual tokens. To understand this misalignment, we first introduce the attention mechanism in the visual encoder and LLM decoder of VLMs in Sec. 3.1 as preliminaries. Then, in Sec. 3.2 and Sec. 3.3, we describe two phenomena observed in the visual attention of LLM decoder, termed attention shift and attention dispersion, corresponding to the position and intensity of attention respectively. These phenomena are absent in the visual encoder of VLMs, which motivates our use of [CLS] attention as a more reliable indicator of visual token importance.
@@ 3.1 Preliminaries
Existing VLMs typically consist of two core components: a visual encoder and an LLM decoder, both are the Transformer-based architecture [56, 12]. Although they both rely on the self-attention mechanism, there are subtle differences in specific implementations. In VLMs, the visual encoder (e.g. CLIP [45]) employs a global attention mechanism, where the patch-wise image ${\bm{X}}=[{{\bm{x}}}_{\text{cls}};{{\bm{x}}}_{\text{img}}^{1},{{\bm{x}}}_{\text{img}}^{2},\cdots,{{\bm{x}}}_{\text{img}}^{n}]\in{\mathbb{R}}^{(n+1)\times d}$ is first converted into the query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ through three weight matrices ${{\bm{W}}}_{Q},{{\bm{W}}}_{K},{{\bm{W}}}_{V}\in{\mathbb{R}}^{d\times d}$ respectively:
@@
While the LLM decoder utilizes causal self-attention [56], where each token can only attend to the previous tokens (i.e. those from the past) and not the following ones (i.e. those from the future). For the whole input sequence ${\bm{X}}=[{{\bm{x}}}_{\text{sys}};{{\bm{x}}}_{\text{img}};{{\bm{x}}}_{\text{txt}};{{\bm{x}}}_{\text{out}}]\in{\mathbb{R}}^{l\times d}$, the query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ are obtained in the same way, and the causal attention is then calculated as:
${\bm{A}}=\text{Softmax}\left(\frac{{\bm{Q}}{{\bm{K}}}^{T}+{\bm{M}}}{\sqrt{d}}\right),\quad{\bm{O}}={\bm{A}}{\bm{V}}$ (3)
where $l$ is the total length of the input sequence (including system prompt, image patches, text question, and output answer), and ${\bm{M}}\in{\mathbb{R}}^{l\times l}$ is a lower triangular causal mask ensuring that each token attends only to itself and previous tokens. Our focus is on the attention received by visual tokens from three sources: image attention (from the image patches themselves), text attention (from the text prompt), and last attention (from the output answer since we only extract attention at the prefill phase, where the output contains only a single token). Multi-head attention is averaged across all heads following FastV [8] for analysis.
@@ 3.2 Text-Visual Attention Shift
Figure 3: Distribution of attention received by visual tokens from different sources, arranged from left to right: [CLS] token, image patches, text prompt, and last token (for output generation). From attention distributions within the LLM decoder (image, text, and last), we observe an obvious trend where attention intensity gradually increases with token indices, a phenomenon we term attention shift. Notably, the same trend does not show up in [CLS] attention.
To analyze text-visual attention in VLMs, we first randomly sample $N$ image-text pairs from the LLaVA-mix665k data [32]. These samples are then used to prompt the VLM to generate responses, during which we extract the attention received by the visual tokens. In this section, $N$ is set to 1,000, and we use LLaVA-1.5-7B as the VLM, following the original configuration in [32]. We analyze the distribution of attention scores with respect to token indices, with sources including the [CLS] token from the visual encoder, as well as image patches, text prompt, and output response from the LLM decoder (after layer 2, following [8]).
As shown in Fig. 3, a clear trend is visible across all attention distributions within the LLM decoder (image, text, and last), where attention scores increase with larger token indices. This suggests that if visual tokens were pruned based on text-visual attention from the LLM decoder, most retained visual tokens would be located in the lower half of the input image, potentially leading to a serious loss of important visual information. We attribute this phenomenon to the unidirectional nature of attention in the LLM, as this trend does not appear in the visual encoder, which employs global attention. Although this causal attention is the core of the next-token prediction paradigm, it is not well-suited for assessing the importance of visual tokens. As shown in the distribution plots, attention from the [CLS] token serves as a more reliable indicator for token pruning. In addition, the red vertical lines in Fig. 3 denote the length of each row in the input image (24 for CLIP-ViT-L-14-336px in LLaVA-1.5). Compared to the attention in the LLM decoder, [CLS] attention is more sensitive to image boundary areas, allocating more attention to central regions containing richer information, which further supports its effectiveness for evaluating visual token importance.

### v2

# Beyond Text-Visual Attention: Exploiting Visual Cues for Effective Token Pruning in VLMs

@@ Abstract
Large vision-language models (LVLMs) generally contain significantly more visual tokens than their textual counterparts, resulting in a considerable computational burden. Recent efforts have been made to tackle this issue by pruning visual tokens early within the language model. Most existing works use attention scores between text and visual tokens to assess the importance of visual tokens. However, in this study, we first analyze the text-visual attention in the language model and find that this score is not an ideal indicator for token pruning. Based on the analysis, We propose VisPruner, a plug-and-play method that utilizes visual cues for more effective token pruning in LVLMs. Specifically, we first use visual attention to select a limited number of significant tokens. Then, we remove duplicate tokens from the remaining ones based on their similarity. By retaining diverse tokens alongside the initially selected important tokens, we maximally preserve the visual information of the input image. Experimental results demonstrate that our VisPruner sustains strong performance across various VLM architectures and reduction ratios, significantly outperforming existing methods based on text-visual attention. Notably, without any training, VisPruner can reduce the FLOPs of LLaVA-1.5-7B by 91% and inference latency by 75%, while maintaining comparable performance. Our code is available at https://github.com/Theia-4869/VisPruner.
@@ 1 Introduction
Figure 1: The illustration of different pruning methods. Correct answer parts are shown in blue, while hallucinations due to pruning are shown in red. Text-visual attention methods like FastV often preserve the lower parts of input images, which can lead to the loss of crucial visual information during early pruning (e.g. the iron in the man’s hand). Random pruning removes positional bias but still fails to preserve important visual information, leading to even more hallucinations (e.g. suitcase and stop sign). Our VisPruner, which uses visual cues for pruning, effectively answers the question with a significant token reduction ratio and provides more details (e.g. the type and color of the car, shown in green).
@@
With recent advances in large language models (LLMs) [42, 50, 24, 2], numerous efforts have been made to extend their powerful reasoning capabilities to multi-modal tasks [15, 37, 61], giving rise to vision language models (VLMs) [1, 48, 36, 3, 10, 54]. Existing VLMs typically contain a visual encoder (e.g. CLIP [43]) to convert visual inputs into sequential representations, a language model (e.g. Llama 2 [51]) to process questions and generate answers, and a modal alignment module (e.g. MLP, Q-Former) for the language model to accept visual tokens as input [36, 3]. Although these approaches have achieved remarkable performance in multi-modal tasks like visual question answering [16, 22, 17, 38, 46], the visual tokens, which comprise the majority of the input sequence, significantly increase the computational cost of VLM inference. For instance, LLaVA-1.5 [34] accepts $336\times 336$ input images, which are converted into 576 visual tokens, while their textual counterparts have typically fewer than 100 tokens. Higher input resolutions or video streams lead to even longer visual token sequences (e.g. 2880 for LLaVA-NeXT [35] and 2560 for Video-LLaVA [32]), further escalating computational load.
To reduce the inference cost of VLMs, FastV [9] initially attempts early pruning of visual tokens within the language model. It identifies that, in the deeper layers of the language model, a substantial number of visual tokens receive significantly less attention than their textual counterparts, and discards many of these low-attention visual tokens after layer 2. Subsequently, abundant recent work follows this paradigm, pruning visual tokens based on text-visual attention from the language model [60, 65, 55]. Although visual tokens in VLM are redundant, we find that text-visual attention within the language model may not be an ideal indicator for token pruning. As illustrated in  Fig. 1, pruning based on text-visual attention suffers from positional bias, primarily preserving the lower parts of the image with less useful information (or even completely ineffective paddings). Meanwhile, tokens in the center region with crucial visual information (e.g. the iron) are discarded, making it difficult for the model to answer correctly.
We conduct an in-depth analysis of text-visual attention in the language model and make two findings: (1) Text-visual attention shift. Affected by the long-term decay property of rotary position embedding in the language model [47], for inputs structured as [image, text], text tokens located later tend to pay higher attention towards visual tokens that are physically closer to them (i.e. patches at the lower part of the input image), a phenomenon that exists from the first layer [18]. (2) Text-visual attention dispersion. Although adjusting the position embedding of the visual token part can avoid the shift issue, both before and after the modification, the same problem persists: the attention density curve exhibits high entropy and low peak, resembling a more uniform distribution, making it difficult to select actually important visual tokens. While these phenomena do not impact the inference process with full visual tokens, pruning based on text-visual attention significantly degrades performance, especially at higher reduction ratios.
Based on our observations, we propose VisPruner, a plug-and-play approach that exploits visual cues for more effective token pruning in VLMs. Our VisPruner first selects a small number of important tokens with rich visual information based on the attention of the visual encoder. Then, to ensure that the retained tokens do not focus solely on the foreground and overlook background information, we eliminate duplicate tokens from the remainder based on similarity. The remaining diverse tokens, along with the initially selected important tokens, maximize the retention of visual information in the input image. Notably, unlike text-visual attention-based methods that prune within the language model, our approach leverages visual cues to prune before the language model, ensuring compatibility with various attention optimization techniques (e.g. FlashAttention [12]) without the need for full attention within the language model, thus achieving higher inference efficiency.
@@
Benefiting from its simple design, we can easily apply VisPruner to different VLMs, including the LLaVA series [34, 35, 32] and other architectures like Qwen-VL [3]. Extensive experiments on various vision-language benchmarks demonstrate that VisPruner consistently outperforms other text-visual attention-based methods under different reduction ratios. For example, VisPruner is capable of pruning 94.4% of visual tokens in LLaVA-1.5-7B, reducing inference FLOPs by more than 95%, and maintaining 91.5% of the original performance across 10 tasks without any additional training. In scenarios involving longer visual token sequences, such as high-resolution images (LLaVA-NeXT) or video inputs (Video-LLaVA), VisPruner continues to achieve superior performance with its simple design. In summary, the contributions of our work are three-fold:
@@
    We conduct a thorough investigation of text-visual attention in the language model and discover that this attention is not an ideal indicator for token pruning.
@@
    We introduce VisPruner, a plug-and-play approach that exploits visual cues for more effective token pruning in VLMs based on our observations.
@@
    We apply VisPruner to various VLMs and conduct extensive experiments across diverse vision-language tasks, demonstrating our method can maintain high performance even under extreme reduction ratios.
@@ 2 Related Work
Vision-language models (VLMs). The recent impressive success of large language models (LLMs) [42, 50, 24, 2] leads to a trend of extending their powerful reasoning capabilities to multi-modal comprehension tasks, giving birth to vision-language models (VLMs) [1, 48]. These VLMs generally include a visual encoder for serializing representations of input images and a language model for text generation. To make the language model acceptable for visual representations as input, they typically align the visual and language modalities with an alignment module, which can be a simple linear layer [36], an MLP projector [34], or a deep query-based network [27, 3]. Although this allows the language model to have visual perception, the introduction of long visual token sequences increases the computational burdens. Furthermore, studies have shown that existing VLMs still suffer from visual shortcomings [49] or hallucinations [20]. To mitigate this, efforts have been made to improve their performance by increasing the resolution of input images [39, 58], which further exacerbates the computational burden. Optimizing the inference process of VLMs is an urgent task to enable their application in resource-constrained real-world scenarios.
Token reduction for VLMs. One way to optimize VLM inference is by reducing the visual tokens that occupy the majority of the input sequence. Several studies have explored token reduction in language models [11, 21, 41, 44]. Compared with text, image information tends to have higher redundancy, making visual token reduction for VLMs more reasonable and effective. Various efforts have been made to compress visual tokens into a more compact representation [45, 8, 31, 64, 28, 19, 5]. However, these methods require additional training, increasing computational burden and training costs. In this work, we mainly focus on training-free VLM inference optimization, typically achieved by removing the less important tokens. FastV [9] first identifies the redundancy and inefficiency of visual tokens in VLMs and proposes a simple method to remove visual tokens with low attention scores after the second layer of the language model. SparseVLM [65] removes distractions from text prompts and uses more accurate text attention to progressively sparsify visual tokens. After this, an increasing number of studies choose to rely on text-visual attention within the language model to assess the importance of visual tokens [60, 55]. However, after an in-depth investigation, we find that such text-visual attention may not be an ideal indicator for token pruning. Unlike these previous work, we propose to exploit visual cues for more effective token pruning. Notably, some concurrent work shares similar conclusions with us [66, 53, 59].
3 Text-Visual Attention Investigation
Most prior research utilizes text-visual attention in the language model to evaluate the significance of visual tokens for pruning. In this section, we conduct an in-depth investigation and suggest that this attention may not be an ideal indicator for token pruning. We begin in Sec. 3.1 by introducing the attention mechanisms in VLMs as preliminaries. Then, in Sec. 3.2 and Sec. 3.3, we analyze two phenomena observed in text-visual attention, termed attention shift and attention dispersion, which are influenced by positional bias and density distribution, respectively.
@@ 3.1 Preliminaries
Existing VLMs generally consist of two core components: a visual encoder and a language model, both are the Transformer-based architecture [52, 14]. Although they both rely on the self-attention mechanism, there are subtle differences in specific implementations. The visual encoder (e.g. CLIP [43]) employs a global attention mechanism, where the patch-wise image ${\bm{X}}=[{{\bm{x}}}_{\text{cls}};{{\bm{x}}}_{\text{img}}^{1},{{\bm{x}}}_{\text{img}}^{2},\cdots,{{\bm{x}}}_{\text{img}}^{n}]\in{\mathbb{R}}^{(n+1)\times d}$ is first converted into the query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ through three weight matrices ${{\bm{W}}}_{Q},{{\bm{W}}}_{K},{{\bm{W}}}_{V}\in{\mathbb{R}}^{d\times d}$ respectively:
@@
While the language model (e.g. Llama 2 [51]) utilizes causal self-attention [52], where each token can only attend to the previous tokens (i.e. those from the past) and not the following ones (i.e. those from the future). For the whole input sequence ${\bm{X}}=[{{\bm{x}}}_{\text{sys}};{{\bm{x}}}_{\text{img}};{{\bm{x}}}_{\text{txt}}]\in{\mathbb{R}}^{l\times d}$, the query ${\bm{Q}}$, key ${\bm{K}}$ and value ${\bm{V}}$ are obtained in the same way, and the causal attention with relative rotary position embedding is then calculated as:
${\bm{A}}=\text{softmax}\left(\frac{({{\bm{R}}}_{\theta}{\bm{Q}}){({{\bm{R}}}_{\theta}{{\bm{K}}})}^{T}+{\bm{M}}}{\sqrt{d}}\right),\ {\bm{O}}={\bm{A}}{\bm{V}}$ (3)
where $l$ is the total length of the input sequence (including system prompt, image patches and text question), ${{\bm{R}}}_{\theta}\in{\mathbb{R}}^{d\times d}$ is the block-diagonal rotation matrix, and ${\bm{M}}\in{\mathbb{R}}^{l\times l}$ is a lower triangular causal mask ensuring that each token attends only to itself and previous tokens. We refer to the attention received by visual tokens received from the last text token as last attention following [9]. And multi-head attention is averaged across different heads for analysis.
@@ 3.2 Text-visual attention shift
Figure 2: Analysis of text-visual attention shift. (a) Selection frequency and received attention of visual tokens. There is a clear positive correlation between selection frequency and the attention received, which is also accompanied by a positional bias. (b) Proportion of visual tokens with top 25% text-visual attention across each quartile position. In the shallower layers of the language model, attention is notably concentrated on visual tokens with larger indices. (c) Performance with only visual tokens from each quartile position. In the shallower layers, retaining visual tokens located in the central region yields higher performance.
To analyze text-visual attention in VLMs, we first randomly sample $N$ image-text pairs from the LLaVA-mix665k data [34]. These samples are then used to prompt the VLM to generate responses, during which we extract the attention received by the visual tokens. In this section, $N$ is set to 1,000, and we use LLaVA-1.5-7B as the VLM.
Following FastV [9], we prune 75% visual tokens after layer 2 of the language model based on text-visual attention and visualize the frequency with which visual tokens at various positions are selected to be retained, as well as their average attention received across $N$ samples. As shown in Fig. 2(a), there is a clear positive correlation between the selection frequency and the attention received, along with a noticeable shift in text-visual attention. Text tokens tend to focus on visual tokens with higher indices, due to the long-term decay property of rotary position embedding [47]. In the input sequence of the language model, text tokens are located after visual tokens, thus text tokens exhibit higher attention towards later-positioned visual tokens (i.e. closer to themselves). We demonstrate the relationship between this shift and the language model layer index in Fig. 2(b). It can be seen that positional bias appears from the first layer and is more pronounced in the shallower layers, aligning with conclusions from previous work [18].
A natural question then arises: In the presence of inherent positional bias, do higher attention scores necessarily represent richer visual information? To answer this question, we separately preserve visual tokens from each quartile position and evaluate the average performance on four benchmarks of GQA [22], TextVQA [46], POPE [30], and MME [15] at different layers, as shown in Fig. 2(c). To our surprise, in the shallower layers of the model, visual tokens located in the central region of the image maintain the highest performance, rather than those positioned later, which received the most text attention. The difference between these two diminishes as the positional bias weakens beyond the tenth model layer. However, optimizing inference efficiency requires us to prune visual tokens as early as possible within the language model, indicating that the phenomenon of text-visual attention shift indeed has a negative impact on performance after pruning.